# Saupfercheckweg 1, 69117 Heidelberg, Germany

### v1

@@ 1 Introduction
Amongst neutrino mass generation mechanisms, the type-I seesaw scenario requires extension in the fermion sector Minkowski:1977sc ; Mohapatra:1979ia ; Yanagida:1979as ; Gell-Mann:1979vob ; Glashow:1979nm . Such BSM fermions may either themselves be “sterile neutrino dark matter” Dodelson:1993je ; Shi:1998km , or constitute the so-called “neutrino-portal” Falkowski:2009yz ; Lindner:2010rr ; Gonzalez-Macias:2016vxy that may mediate between DM and the SM. However, a purely scalar extension of the SM can give rise to neutrino masses just as well. The simplest instance is the type-II seesaw scenario Schechter:1980gr ; Cheng:1980qt ; Mohapatra:1980yp ; Lazarides:1980nt , in which an electroweak triplet scalar is added to the SM. Its most distinctive feature, perhaps, is the existence of a doubly charged scalar in the spectrum. Doubly charged scalars can also participate in neutrino mass generation in more complex or loop-suppressed scenarios. Such additional scalars can act as portals mediating between scalar DM and the SM thermal bath in the early Universe through the scalar potential. Such scalar-neutrino-portal DM scenarios often lead to leptophilic DM, which is less affected by the DMDD constraints, due to loop suppressed interactions Barman:2021hhg . Importantly, the annihilation products of such DM feature a specific flavor structure that is linked to neutrino oscillation observables, as schematically depicted in Fig. 1.
@@ 2 Neutrino mass models and charged-scalar-portal dark matter
Many theories propose that neutrinos, assumed to be Majorana particles, obtain their masses from higher-dimensional operators; the lowest being the dimension-five operator $\mathcal{O}_{1}=L^{i}L^{j}H^{k}H^{l}\epsilon_{ik}\epsilon_{jl}/\Lambda$, suppressed by a high-energy scale ($\Lambda$) linked to lepton number violating BSM physics. In the above, $L$ represents the lepton doublet and $H$ the SM Higgs doublet with the $SU(2)_{L}$ indices $i,j,k,l$. The seesaw mechanism exemplifies such kind of BSM scenarios. Phenomenologically, greater interest and testability arises when the scale of lepton number violation lies proximate to the electroweak scale, typically within the sub-TeV or TeV range, so that it can be probed in collider and other experiments. The required suppression can be attributed to loop factors, higher-dimensional operators, or a combination thereof. While many neutrino mass models rely on matter field extensions and the incorporation of right-handed neutrinos, our focus is directed toward models where minimal scalar field extensions alone suffice to generate neutrino masses and mixings. Specifically, we investigate scenarios wherein a doubly charged scalar emerges.
@@
On the other hand, BSM particles responsible for generating neutrino mass could potentially connect to DM, offering intriguing new physics that establishes flavor-specific DM indirect signatures directly correlated with neutrino oscillation data. Our work investigates the role of such scalars introduced for neutrino mass generation in DM phenomenology. We focus on a minimal scalar DM candidate stabilized by a discrete $\mathbb{Z}_{2}$ symmetry. Generally, the DMDD constraints restrict this scenario to DM masses in the vicinity of Higgs resonance $m_{\mathrm{DM}}\sim m_{h}/2$ or in a narrow range around $3\ \mathrm{TeV}$ Cirelli:2024ssz . The scalars responsible for neutrino mass generation, however, couple to the scalar DM (we call it ‘scalar-neutrino-portal’) and provide additional annihilation channels that are not directly connected to DMDD, thus reducing the constraints significantly. On the other hand, the doubly charged scalars could predominantly couple to leptons and provide flavour specific signals both in collider and indirect search experiments, shedding light on neutrino properties like mass ordering and Majorana phases.
@@ 3 Scalar-neutrino-portal in the type-II seesaw model
Our focus here will be to elaborate on the scalar-neutrino-portal in seesaw models in which neutrino masses are generated at tree-level. As a specific case study, we scrutinize the type-II seesaw framework, wherein the scalar sector is expanded to include an $SU(2)_{L}$ scalar triplet $\Delta$ with a hypercharge $Y=+1$ alongside a complex scalar singlet $S$, which undergoes $S\rightarrow-S$ under a discrete $\mathbb{Z}_{2}$ symmetry and accounts for DM. [^2] A real scalar works equally, complex scalar is just a choice with an additional degree of freedom. The SM fermionic and gauge sector remains unaltered. The general renormalizable scalar potential reads
@@
Before analyzing interconnections with DM, it is important to emphasize that the SM predicts the absence of lepton flavor-violating processes. However, an inevitable expectation of lepton flavor violation arises as neutrino mass models emerge to explain oscillation data by generating neutrino masses and mixings. Various neutrino mass models yield different predictions of lepton flavor-violating observables. Notably, within the type-II seesaw mechanism, thorough analyses of these observables have been conducted Primulando:2019evb ; Fridell:2023gjx . Here we present a brief summary of these lepton flavor-violating observables and the constraints placed on the Yukawa matrix $y_{L\Delta}$ since the Yukawa interactions induce lepton flavor-violating phenomena, such as $\ell_{a}^{-}\rightarrow\ell_{b}^{+}\ell_{c}^{-}\ell_{d}^{-}$ and $\ell_{a}^{-}\rightarrow\ell_{b}^{-}\gamma$, which are tightly constrained (cf. Tab. 2).
@@
For simplicity, we consider the scenario where $\Delta^{++}$ is the lightest of the fields from the triplet. Now, the doubly charged scalar decay modes depend on the vev $v_{\Delta}$ and mass splitting between the fields from the triplet, and this can be nicely presented in decay phase diagrams as shown in Ref. Melfo:2011nx . Since we are concentrating on leptonic decay modes of the doubly charged scalar, we set $v_{\Delta}\ll 10^{-4}\ \mathrm{GeV}$ such that the leptonic branching fraction is $100\%$ and we set the mass splitting very small such that cascade decays do not open up. The decay rate of the doubly charged scalar $\Delta^{++}$ into a pair of same-sign charged leptons can be expressed as Akeroyd:2007zv ; Melfo:2011nx ; Mandal:2022zmy
@@ 5 Dark Matter phenomenology
As mentioned above, we consider DM to be a complex scalar in this work, stabilized by a discrete $\mathbb{Z}_{2}$ symmetry that prevents its decay. We focus on the WIMP regime, where the coupling of DM to the SM is strong enough to ensure thermalisation in the early Universe. Annihilation processes keep the DM abundance to its equilibrium abundance as the bath temperature $T$ drops below the DM mass and the DM abundance gets Boltzmann suppressed. The relic abundance is finally determined by the freeze-out of DM annihilation processes, i.e. the time when the annihilation rate drops below the Hubble rate. The WIMP relic abundance is inversely proportional to the thermally averaged annihilation cross section $\braket{\sigma v}$. To reproduce the observed relic abundance $(\Omega_{\mathrm{DM}}h^{2})_{\mathrm{obs}}=0.1200\pm 0.0012$ Planck:2018vyg , we require (e.g. Steigman:2012nb )
@@ 6 Dark matter indirect detection and flavor dependence
where $D(E,\vec{x})$ is a diffusion coefficient, $b(E,\vec{x})$ the effective energy loss rate444 It has been pointed out recently that the discontinuous energy loss of multi-TeV positrons experiencing inverse Compton scattering can enhance the dark matter signal close to the endpoint John:2023ulx (see also Ref. Zdziarski1989ApJ…342.1108Z ) compared to the continuous approximation. For the energies considered here, the enhancement is expected to be $\mathcal{O}(0.1)$. and $Q(E,\vec{x})$ the cosmic ray source term. The source term due to DM annihilation is given by
@@ 6.2 Annihilation in the smooth Milky Way halo
Fig. 9 shows the expected positron fluxes for two different DM/mediator mass benchmarks and different neutrino parameters. Compared to the injection spectra in Fig. 8, the peaks are smoothed out, but the drop towards their endpoints still carries some flavor information. Concretely, the larger the DM annihilation fraction into electrons, the more pronounced is the drop-off of the spectra towards the kinematic endpoint. This may in principle enable e.g. distinguishing the type-II seesaw IH scenario from the NH or Zee-Babu scenarios. However, the overall flux is low – more than an order of magnitude below current AMS-02 data AMS:2021nhj ; Maurin:2013lwa and possibly beyond the reach of AMS-100 Schael:2019lvx for TeV-scale DM masses.
As the annihilation rate scales as $\rho^{2}$, there can be significant enhancement by DM substructure on scales smaller than the positron diffusion scale. Such an astrophysical “boost factor” is expected in the present WIMP DM scenario and can lead to orders-of-magnitude larger cosmic ray fluxes (e.g. Refs. Delos:2022bhp ; Stucker:2023rjr ), bringing the expected flux closer to future detectability. The overall emission morphology in the Milky Way would however be similarly extended as in the smooth-halo case shown here, and we hence expect similar spectral shapes of the resulting fluxes.
@@ 6.3 Signal from a nearby subhalo
Fig. 9 shows the expected positron fluxes from a nearby subhalo as dashed lines for the same DM and neutrino benchmarks as considered in the previous subsection. Depending on the produced primary lepton flavor composition, the spectra are sharply peaked and the different neutrino scenarios result in distinctive spectral shapes. The normalisation of the spectra depends linearly on the subhalo luminosity and distance. For $m_{\mathrm{DM}}=1\ \mathrm{TeV}$, the benchmark subhalo would be clearly discernible in present data, while a less luminous subhalo or a more massive DM candidate may be detected by future experiments.
@@ 6.4 Discussion
Without a boost from substructure, the positron flux from DM annihilation in the present scenario is too small to be detected in present data, or with a more powerful future experiment. However, small-scale substructure in DM microhalos Delos:2022bhp ; Stucker:2023rjr may enhance the generic WIMP expectation for the Galactic DM annihilation rate by an order of magnitude, which could bring the Galactic signal closer to detectability.
@@ 7 Collider complementarity
Figure 10: Sensitivities at HL-LHC for neutrino oscillation parameters with respect to the neutrino mass scale $m_{0}$ and the Majorana phase $\varphi$ (we assume $\varphi_{1}=\varphi_{2}\equiv\varphi$ and the Dirac CP phase from Tab. 1) in the type-II seesaw model for a fixed mediator mass $m_{\Delta}=1\ \mathrm{TeV}$. The contours indicate the statistical significance at $3\,\mathrm{ab}^{-1}$ integrated luminosity and the red, dashed lines show the $3\sigma$ contours for different integrated luminosities.

### v2

@@ 1 Introduction
Amongst neutrino mass generation mechanisms, the type-I seesaw scenario requires extension in the fermion sector Minkowski:1977sc ; Mohapatra:1979ia ; Yanagida:1979as ; Gell-Mann:1979vob ; Glashow:1979nm . Such BSM fermions may either themselves be “sterile neutrino dark matter” Dodelson:1993je ; Shi:1998km , or constitute the so-called “neutrino-portal” Falkowski:2009yz ; Lindner:2010rr ; Gonzalez-Macias:2016vxy that may mediate between DM and the SM. However, a purely scalar extension of the SM can give rise to neutrino masses just as well. The simplest instance is the type-II seesaw scenario Schechter:1980gr ; Cheng:1980qt ; Mohapatra:1980yp ; Lazarides:1980nt , in which an electroweak triplet scalar is added to the SM. Its most distinctive feature, perhaps, is the existence of a doubly charged scalar in the spectrum. Doubly charged scalars can also participate in neutrino mass generation in more complex or loop-suppressed scenarios. Such additional scalars can act as portals mediating between scalar DM and the SM thermal bath in the early Universe through the scalar potential. Such scalar-neutrino-portal DM scenarios often lead to leptophilic DM, which is less affected by the DMDD constraints, due to loop suppressed interactions Barman:2021hhg . Importantly, the annihilation products [^2] The dark matter candidate can annihilate into neutrinos, but this channel is less relevant for detection, as it also annihilates into charged leptons. Notable sensitivity differences between these channels are highlighted in Ref. Cirelli:2024ssz . of such DM feature a specific flavor structure that is linked to neutrino oscillation observables, as schematically depicted in Fig. 1.
@@
In DM indirect detection (DMID), the couplings of the mediator to the lepton flavors determine the spectrum of cosmic ray positrons expected from DM annihilation in our Galaxy. This has received attention in the light of various cosmic ray excesses observed over the past decades (see in particular Refs. Gogoladze:2009gi ; Dev:2013hka ; Li:2017tmd ; Li:2018abw ; Sui:2017qra ). Notably, the rise of positron/electron flux ratio towards TeV energies was first observed by PAMELA Adriani:2008zq and ATIC Chang:2008aa (which may plausibly be of astrophysical origin, while DM explanations are subject to strong gamma ray constraints, see e.g. Cirelli:2024ssz ; Lin:2014vja ; Xiang:2017jou ), while a more recent but inconclusive excess in the electron plus positron spectrum has been observed by DAMPE DAMPE:2017fbg .
@@ 2 Neutrino mass models and charged-scalar-portal dark matter
Many theories propose that neutrinos, assumed to be Majorana particles, obtain their masses from higher-dimensional operators; the lowest being the dimension-five operator $\mathcal{O}_{1}=L^{i}L^{j}H^{k}H^{l}\epsilon_{ik}\epsilon_{jl}/\Lambda$, suppressed by a high-energy scale ($\Lambda$) linked to lepton number violating BSM physics Weinberg:1979sa . In the above, $L$ represents the lepton doublet and $H$ the SM Higgs doublet with the $SU(2)_{L}$ indices $i,j,k,l$. The seesaw mechanism exemplifies such kind of BSM scenarios. Phenomenologically, greater interest and testability arises when the scale of lepton number violation lies proximate to the electroweak scale, typically within the sub-TeV or TeV range, so that it can be probed in collider and other experiments. The required suppression can be attributed to loop factors, higher-dimensional operators, or a combination thereof. While many neutrino mass models rely on matter field extensions and the incorporation of right-handed neutrinos, our focus is directed toward models where minimal scalar field extensions alone suffice to generate neutrino masses and mixings. Specifically, we investigate scenarios wherein a doubly charged scalar emerges.
@@
On the other hand, BSM particles responsible for generating neutrino mass could potentially connect to DM, offering intriguing new physics that establishes flavor-specific DM indirect signatures directly correlated with neutrino oscillation data. Our work investigates the role of such scalars introduced for neutrino mass generation in DM phenomenology. We focus on a minimal scalar DM candidate stabilized by a discrete $\mathbb{Z}_{2}$ symmetry. Generally, the DMDD constraints restrict this scenario to DM masses in the vicinity of Higgs resonance $m_{\mathrm{DM}}\sim m_{h}/2$ or in a narrow range around $3\ \mathrm{TeV}$ Cirelli:2024ssz . The scalars responsible for neutrino mass generation, however, couple to the scalar DM (we call it ‘scalar-neutrino-portal’) and provide additional annihilation channels that are not directly connected to DMDD, thus reducing the constraints significantly. On the other hand, the doubly charged scalars could predominantly couple to leptons and provide flavor specific signals both in collider and indirect search experiments, shedding light on neutrino properties like mass ordering and Majorana phases.
@@ 3 Scalar-neutrino-portal in the type-II seesaw model
Our focus here will be to elaborate on the scalar-neutrino-portal in seesaw models in which neutrino masses are generated at tree-level. As a specific case study, we scrutinize the type-II seesaw framework, wherein the scalar sector is expanded to include an $SU(2)_{L}$ scalar triplet $\Delta$ with a hypercharge $Y=+1$ alongside a complex scalar singlet $S$, which undergoes $S\rightarrow-S$ under a discrete $\mathbb{Z}_{2}$ symmetry and accounts for DM. [^3] A real scalar works equally, complex scalar is just a choice with an additional degree of freedom. The SM fermionic and gauge sectors remain unaltered. The general renormalizable scalar potential reads
@@
Before analyzing interconnections with DM, it is important to emphasize that the SM predicts the absence of lepton flavor-violating processes. However, an inevitable expectation of lepton flavor violation arises as neutrino mass models emerge to explain oscillation data by generating neutrino masses and mixings. Various neutrino mass models yield different predictions of lepton flavor-violating observables. Notably, within the type-II seesaw mechanism, thorough analyses of these observables have been conducted BhupalDev:2018tox ; Primulando:2019evb ; Fridell:2023gjx . Here we present a brief summary of these lepton flavor-violating observables and the constraints placed on the Yukawa matrix $y_{L\Delta}$ since the Yukawa interactions induce lepton flavor-violating phenomena, such as $\ell_{a}^{-}\rightarrow\ell_{b}^{+}\ell_{c}^{-}\ell_{d}^{-}$ and $\ell_{a}^{-}\rightarrow\ell_{b}^{-}\gamma$, which are tightly constrained (cf. Tab. 2).
@@
For simplicity, we consider the scenario where $\Delta^{++}$ is the lightest of the fields from the triplet. Now, the doubly charged scalar decay modes depend on the vev $v_{\Delta}$ and mass splitting between the fields from the triplet, and this can be nicely presented in decay phase diagrams as shown in Ref. Melfo:2011nx . Since we are concentrating on leptonic decay modes of the doubly charged scalar, we set $v_{\Delta}\ll 10^{-4}\ \mathrm{GeV}$ such that the leptonic branching fraction is $100\%$ and we set the mass splitting very small such that cascade decays do not open up. The decay rate of the doubly charged scalar $\Delta^{++}$ into a pair of same-sign charged leptons can be expressed as Akeroyd:2007zv ; FileviezPerez:2008jbu ; Melfo:2011nx ; Mandal:2022zmy
@@ 5 Dark Matter phenomenology
As mentioned above, we consider DM to be a complex scalar in this work, stabilized by a discrete $\mathbb{Z}_{2}$ symmetry that prevents its decay. We focus on the WIMP regime, where the coupling of DM to the SM is strong enough to ensure thermalisation in the early Universe. Annihilation processes keep the DM abundance to its equilibrium abundance as the bath temperature $T$ drops below the DM mass and the DM abundance gets Boltzmann suppressed. The relic abundance is finally determined by the freeze-out of DM annihilation processes, i.e. the time when the annihilation rate drops below the Hubble rate. The WIMP relic abundance is inversely proportional to the thermally averaged annihilation cross section $\braket{\sigma v}$. To reproduce the observed relic abundance $(\Omega_{\mathrm{DM}}h^{2})_{\mathrm{obs}}=0.1200\pm 0.0012$ Planck:2018vyg , we require (e.g. Steigman:2012nb ; Bringmann:2020mgx )
@@ 6 Dark matter indirect detection and flavor dependence
where $D(E,\vec{x})$ is a diffusion coefficient, $b(E,\vec{x})$ the effective energy loss rate555 It has been pointed out recently that the discontinuous energy loss of multi-TeV positrons experiencing inverse Compton scattering can enhance the dark matter signal close to the endpoint John:2023ulx (see also Ref. Zdziarski1989ApJ…342.1108Z ) compared to the continuous approximation. For the energies considered here, the enhancement is expected to be $\mathcal{O}(0.1)$. and $Q(E,\vec{x})$ the cosmic ray source term. The source term due to DM annihilation is given by
@@ 6.2 Annihilation in the smooth Milky Way halo
Fig. 9 shows the expected positron fluxes for two different DM/mediator mass benchmarks and different neutrino parameters. Compared to the injection spectra in Fig. 8, the peaks are smoothed out, but the drop towards their endpoints still carries some flavor information. Concretely, the larger the DM annihilation fraction into electrons, the more pronounced is the drop-off of the spectra towards the kinematic endpoint. This may in principle enable e.g. distinguishing the type-II seesaw IH scenario from the NH or Zee-Babu scenarios, assuming the astrophysical background is sufficiently smooth. However, the overall flux is low – more than an order of magnitude below current AMS-02 data AMS:2021nhj ; Maurin:2013lwa and possibly beyond the reach of AMS-100 Schael:2019lvx for TeV-scale DM masses.
As the annihilation rate scales as $\rho^{2}$, there can be significant enhancement by DM substructure on scales smaller than the positron diffusion scale. Such an astrophysical “boost factor” is expected in the present WIMP DM scenario and can lead to 1-2 orders-of-magnitude larger cosmic ray fluxes (e.g. Refs. Delos:2022bhp ; Stucker:2023rjr ), bringing the expected flux closer to future detectability. Even larger boost factors have been invoked in the literature, but are subject to stringent gamma-ray constraints (e.g. Lin:2014vja ).
The overall emission morphology in the Milky Way would however be similarly extended as in the smooth-halo case shown here, and we hence expect similar spectral shapes of the resulting fluxes.
@@ 6.3 Signal from a nearby subhalo
Fig. 9 shows the expected positron fluxes from a nearby subhalo as dashed lines for the same DM and neutrino benchmarks as considered in the previous subsection. Depending on the produced primary lepton flavor composition, the spectra are sharply peaked and the different neutrino scenarios result in distinctive spectral shapes. The normalisation of the spectra depends linearly on the subhalo luminosity and distance. For $m_{\mathrm{DM}}=1\ \mathrm{TeV}$, the benchmark subhalo could have already shown up in present data, while a less luminous subhalo may be detected by future experiments. For $m_{\mathrm{DM}}=10\ \mathrm{TeV}$, the benchmark subhalo could generate a clear signal at future experiments, with the potential to discriminate between the different neutrino scenarios.
@@ 6.4 Discussion
Without a boost from substructure, the positron flux from DM annihilation in the present scenario is too small to be detected in present data, or with a more powerful future experiment. However, small-scale substructure in DM microhalos Ishiyama:2010es ; Delos:2022bhp ; Stucker:2023rjr may enhance the generic WIMP expectation for the Galactic DM annihilation rate by an order of magnitude, which could bring the Galactic signal closer to detectability.
@@
Note that the identification of DM-induced features in the positron flux depends on the smoothness or subdominance of astrophysical backgrounds. While present data indicate a smooth flux amenable to DM analyses John:2021ugy , effects of distinct nearby positron sources may result in spectral irregularities Cholis:2021kqk ; Krommydas:2022loe ; Lv:2024khg (see however John:2022asa ). This would hamper the inference of neutrino properties when the DM-induced flux does not clearly dominate over backgrounds.
@@ 7 Collider complementarity
Figure 10: Sensitivities at HL-LHC for neutrino oscillation parameters with respect to the neutrino mass scale $m_{1,3}$ and the Majorana phase $\varphi$ (we assume $\varphi_{1}=\varphi_{2}\equiv\varphi$ and the Dirac CP phase from Tab. 1) in the type-II seesaw model for a fixed mediator mass $m_{\Delta}=1\ \mathrm{TeV}$. The contours indicate the statistical significance at $3\,\mathrm{ab}^{-1}$ integrated luminosity and the red, dashed lines show the $3\sigma$ contours for different integrated luminosities. The straight vertical line indicates the upper bound for the lightest neutrino mass.